1.How many mililiters of a 20% acid solution should Annie mix with a 60% acid solution to make 100 mililiters of a 50% acid solution?


Smeagol kicked a ball. The function h(t)= -16t2 +80t describes the height of the ball, in feet, t seconds
after he kicked it. How many seconds did the ball take to first reach 64 feet?

Answer by Mathtut(3670)

let x and y be the amount of 20% and 60% solutions
:
x+y=100......eq 1
.2x+.6y=.5(100)...eq 2
:
rewrite eq 1 to x=100-y and plug it into eq 2
:
.2(100-y)+.6y=50
:
20-.2y+.6y=50
:
.4y=30
:
y=75 mililiters of 60% solution
:
x=100-y=100-75=25 milliliters of 20% solution